OpenStudy (anonymous):

simplify using identities so that you can solve the equation. Give the general solution in radians. sin x = 1 - 2 cos^2 x

OpenStudy (tkhunny):

Have you considered \(\cos^{2}(x) = 1 - \sin^2(x)\)?

OpenStudy (anonymous):

but what do I do to get the answer in radians?

OpenStudy (trojanpoem):

Do as tkhunny said then Factorize.... sinx = 1- 2 (1-sin^2(x)) sinx = 1 - 2 +2sin^2(x) 2sin^2(x) - sinx - 1 = 0 (2sinx + 1)(sinx - 1) = 0 sinx = 1 sinx = -0.5

OpenStudy (math&ing001):

To get the answer in radians you just put in pi/2 instead of 90° for sinx=1 for example. Or you set you calculator to radians when doing x=sin^-1(-1/2)

OpenStudy (anonymous):

thank you!

OpenStudy (tkhunny):

You simply solve the equation. Why would it be in anything but Radians? BTW "Radians" aren't really a unit of anything. They are just numbers. \(\sin(x) = 1 \implies x = \dfrac{\pi}{2} + 2k\pi\; for\; k\;an\;integer.\) There are infinitely many such solutions.